Find the exponential decay function for a radioactive material that has a half-life of 1234 years. How long will it take until only 10% of the material remains?
I'm just having a problem setting up this problem. Any help you can give will be greatly appreciated.

Answer by stanbon(48558)

Formula: A=P(1/2)^(t/1234)
A is the amount you have after some years.
P is the amount you started with.
t is the number of years since the start time.
You want A=10%(p)= 0.1P
So, you have 0.1P=P(1/2)^(t/1234)
0.1=2^(-t/1234)
Take the ln of both sides to get:
ln(0.1)= (-t/1234)(ln 2)
t=[-1234{ln 0.1)/(ln 2)]
I don't have a calculator with me so I will leave
the calculating to you.
Cheers,
Stan H.